ECD signal stuck at 844883.8

Hi guys,

So I was walking by one of our less frequently used 8890s and saw that all my flow and temperature zones were off and the detectors were saturated at 850k. I'm the only person who operates it and I didn't leave it like that. Checked the system log and it was fine so I started up carrier and aux gases, only problem was the front detector signal didn't move. I tried restarting the instrument, restarting openlab, turned the flow on and off for the aux gas, and pushing large volumes through both the carrier and aux lines.

Is this thing toast? It's only 14 months old.

8/23/2021 edit: I checked the exhaust flows and everything was where it should be so I heated the detectors and made a solvent injection. Back channel looked fine but the front channel never moved.

  • Hi

    Can you confirm, there is only detector or two detector...

  • It is a dual ECD. I'm guessing it's some kind of electronic problem because the number isn't fluctuating at all, it's just sitting at 844883.8. It was at 113 Hz last time I was running it around 2 weeks ago.

  • The only time my ECD was at the max was when the nitrogen flow was zero. Check your flow coming out the vent tube to make sure you have flow at about what it is set at. We did have the flow controller go bad once where it wasn't opening to what it should be.  You will probably need a service call if the flow is ok. Also no it should not be toast after a year.

  • Yeah, I haven’t hooked it up to a flow meter, but I was programming different flows and the instrument was responding to the set points. Even at saturation there should be some variance in the signal, it didn’t change at all for 4 hours. I’m gonna call Agilent on Monday. 

  • Pulled the electronics cover and found a loose connection that goes to the front electrometer. The black ribbon cable on top was only locked in on one side, turned the instrument on and am still getting the same reading. Is there a battery on this board somewhere that's still holding onto settings from when that cable was still loose?

                                           

  • Fixed, the other end of the electrometer ribbon was also barely plugged in. Took the top cover off and plugged the connector in all the way. Signal is back to normal.

Was this helpful?